OpenStudy (anonymous):

If tan x = .67 and cos x = .83, what is sin x?

OpenStudy (tyteen4a03):

Hint: Remember that tan x = sin x / cos x.

hero (hero):

Also, write it as \[\tan \theta = \frac{67}{100}\] and \[\cos \theta = \frac{83}{100}\] And use \(\theta\) , not \(x\) for angles Now use right triangle trigonometry:

OpenStudy (anonymous):

tan = sin / cos sin = tan * cos

OpenStudy (tyteen4a03):

@Hero It's standard to call the angle as theta, but it's really just personal preferences.

hero (hero):

Yeah but that's why students get confused because the unknown side of the triangle is usually labeled \(x\)

OpenStudy (anonymous):

@Hero tyteen is right

hero (hero):

I don't need anyone to tell me who is right. I'm also right as well. It's called "perspective".

OpenStudy (anonymous):

x is written mostly when it is in radians

hero (hero):

It's not my fault that they decided to use x to represent such quantities. "The mathematicians" with all of the different symbols they use could have thought of a symbol besides x to represent radians